# Proposal for a 3-Wave Mixing Element with Quantum Paraelectric Materials

### v1

@@ Abstract
In Fig. 3 we plot $R(\omega)$ for both STO and KTO capacitors, choosing the DC bias voltage $v_{0}=v_{\mathrm{0,max}}$ to maximize $\xi$ for each design. For this operating point we simulate a center frequency and 3-dB bandwidth of $\omega_{0}/2\pi=2.072\,\mathrm{GHz}$ and $\kappa/2\pi=24\,\mathrm{MHz}$ for the STO design, respectively, and $\omega_{0}/2\pi=4.882\,\mathrm{GHz}$ and $\kappa/2\pi=49.5\,\mathrm{MHz}$ for the KTO design, respectively. The operating frequency can be tuned in-situ by changing the DC bias. It can also be modified by designing different capacitor dimensions or, a different inductance.
@@
The compression power of a degenerate parametric amplifier is a critical parameter that determines the maximum gain and operating range of the amplifier. It is set by a combination of effects including Kerr nonlinearity, higher order terms in the circuit Hamiltonian, and pump depletion; as such it is complicated to precisely model. Inductive nonlinearities, e.g. kinetic inductance of the electrodes and the proposed stepped-impedance coupler, may also affect power compression. We begin by estimating power compression due to the Kerr effect. A PANDA with an effective Kerr term of order $K_{\mathrm{eff}}/2\pi\approx 0.1\,\mathrm{Hz}$ and linewidth of $\kappa/2\pi\approx 20\,\mathrm{MHz}$ will require $N\approx 10^{8}$ circulating photons to shift the resonator frequency by a linewidth. At a 2 GHz resonant frequency, this corresponds to a circulating power of $P_{\mathrm{circ}}=N\hbar\omega_{0}\kappa\approx-64\,\mathrm{dBm}$, similar to state-of-the-art KIPAs (e.g. 1-dB compression powers of -50 dBm in Ref. [21], -65 dBm in Ref. [24]), and far greater than JJ-based parametric amplifiers engineered for high power compression (e.g. approximately -100 dBm in Ref. [6]). A three-wave mixing element based on a nanoscale capacitor made from STO or KTO can therefore be expected to have power handling competitive with respect to superconducting alternatives, along with advantages from a compact form factor. To optimize power handling it is important to use thick electrodes to reduce kinetic inductance, and to use high quality materials with large dielectric susceptibility and low loss.
@@
Conclusion — We propose the development of a parametric mixing element (PANDA) based on the tunable permittivity of the quantum paraelectric perovskite materials STO and KTO. First, we derive an interaction Hamiltonian for a parametrically modulated resonator comprised of a nanofabricated parallel plate capacitor made from STO or KTO. We calculate voltage-tunable third- and fourth-order nonlinear terms that allow for three-wave or four-wave mixing dynamics, respectively, in analogy with superconducting parametric amplifiers based on the Josephson effect or kinetic inductance. We predict that the third-order (fourth-order) terms can have order MHz (sub-Hz) magnitude, a large ratio which suggests promise as a degenerate parametric amplifier with favorable dynamic range compared to superconducting alternatives. Due to the high permittivity of STO and KTO, a PANDA can also have a compact layout even at sub-GHz frequencies. We also note that with related materials or modifications from isotope exchange, strain, doping can lead to even stronger nonlinearities and lower loss near zero temperature than modeled here [43].
@@
While we focus specifically on a simple parametric amplifier circuit, we expect the large and tunable permittivity of STO and KTO to be widely useful in other RF/microwave frequency quantum circuits including filters, switches/modulators, non-reciprocal devices, and qubits. Unlike superconducting nonlinearities, STO and KTO are, in-principle, robust to optical illumination and magnetic fields [47, 48]; PANDAs can thus be of use for applications such as microwave-to-optical transduction or axion dark matter detection, where robust parametric mixing elements are desired. This proposal highlights the need for materials and nanofabrication development of STO and KTO microwave devices, and the unexplored application of these devices toward compact and scalable quantum technologies.
@@
Acknowledgments.— The work in the Vučković group has been supported by the Department of Energy under the Q-NEXT program and by a Google AI Research grant. C.S.W. acknowledges the support of the Grainger Fellowship from the University of Chicago. G.S. acknowledges support from the Stanford Bloch Postdoctoral Fellowship. C.P.A is supported by AFOSR FA9550-25-1-0309. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. We thank Christian Haffner, Nathaniel Kinsey, Andrew Higginbotham, Osmond Wen and Nicholas Frattini for helpful discussions.
@@ Appendix A Dielectric Properties
The dielectric properties STO and KTO are well described by a modified Landau-Ginzburg-Devonshire (LGD) theory [37]. In this model, the permittivity is inversely proportional to the curvature of the Gibbs free energy as a function of atomic displacement [49, 50, 51]. At room temperature, both STO and KTO are in a paraelectric phase, where the net atomic displacement is zero. Both materials approach a phase transition at cryogenic temperatures. However, this transition is suppressed by quantum fluctuations, resulting a quantum paraelectric state. In these systems, the free energy is described by two potential wells separated by a barrier, and which correspond to two directions of atomic displacement, Fig. 4a. At low temperatures, the energy barrier is weak enough compared to quantum fluctuations such that the atom occupies a superposition of both wells [33, 34].
Figure 4: At cryogenic temperature the perovskites STO and KTO have multiple low-energy states associated with positive or negative displacement of the central atom [36, 37]. This leads to a quantum paraelectric phase in which the quantum fluctuations (horizontal white arrows) are significant compared to the energy barrier between these two potential wells. Permittivity is proportional to the inverse curvature of the free energy compared to relative displacement. An electric field, $E$, tilts the potential, which constraints the magnitude of fluctuations and reduces permittivity. STO and KTO differ in the depth of the double well potential and its susceptibility to an applied field [50]. (b) Relative permittivity $\varepsilon_{r}=\varepsilon/\varepsilon_{0}$ and (c) loss tangent $\mathrm{tan}(\delta)$, modeled vs. $E$. Calculations use Eq. 4 which is based on the model and material parameters in Ref. [35].
Following work and notation by Vendik et al. [52], the dielectric constant of both STO and KTO are modeled using the modified LGD theory as:
@@
[…]
According to Ref. [35], the loss term $\Gamma(\lambda)=\Gamma_{1}(\lambda)+\Gamma_{2}(\lambda)+\Gamma_{3}$ in Eq. 4 is a combination of the following. (1) Loss related to multi-phonon scattering, which scales with temperature such that $\Gamma_{1}(\lambda)=A_{1}(T/T_{c})^{2}G(\lambda)^{1/2}$, where $A_{1}$ is a constant. (2) Residual piezoelectricity, such that $\Gamma_{2}(\lambda)=A_{2}y(\lambda)$ where $A_{2}$ is a constant and
@@
is a residual ferroelectric displacement. Finally, (3) Piezoelectricity due to charged defects such that $\Gamma_{3}=A_{3}n_{d}$, where $n_{d}$ is the density of charged defects and $A_{3}$ is another proportionality constant. The overall loss tangent is therefore modeled as:
@@
These dielectric properties are relevant at RF and microwave frequency regimes and lower (e.g. characterized at 50 MHz in Ref. [38], 1 GHz in Ref. [41], and 3 GHz in Ref. [53]). At frequencies of order THz and higher, the permittivity decreases due to oscillations associated with optical phonon modes in the crystal, and at optical frequencies STO has a refractive index of only 2.4 [43, 42]. To the author’s knowledge, the precise transition between these two limits has not yet been experimentally characterized but is predicted to be above 100’s of GHz [38, 39].
@@
Capacitance is proportional to permittivity, which in this case is voltage-tunable.
@@ Appendix C Circuit Hamiltonian
where the capacitor’s energy is modified from the usual quadratic to account for a nonlinear capacitance (i.e. a nonlinear charge/voltage relation). This Hamiltonian is comprised of both a linear component $H_{\mathrm{harm}}$ and nonlinear interaction terms $H_{\mathrm{int}}$, such that $H=H_{\mathrm{harm}}+H_{\mathrm{int}}$.
@@
It is often helpful to express Eq. 17 in terms of the quadrature variables $a=\frac{1}{\sqrt{2\hbar}}\left(\frac{\phi}{\sqrt{z_{0}}}+i\sqrt{z_{0}}q_{\delta}\right)$ and $a^{\dagger}=\frac{1}{\sqrt{2\hbar}}\left(\frac{\phi}{\sqrt{z_{0}}}-i\sqrt{z_{0}}q_{\delta}\right)$ such that:
@@
where $\phi_{\mathrm{zpf}}=\sqrt{z_{0}\hbar/2}$ and $q_{\mathrm{zpf}}=\sqrt{\hbar/2z_{0}}$ are the zero point fluctuations of branch flux and charge, respectively [54]. Although for now we treat $\phi$ and $q_{\delta}$ as classical fields, in the quantum mechanical picture these become operators with the commutation relations $\hat{a},\hat{a}^{\dagger}]=1$ and $\hat{\phi},\hat{q}_{\delta}]=i\hbar$. In general these variables depend on bias charge since $z_{0}(v)=\sqrt{L/C(v)}$.
We express the quadrature operators in terms of the differential charge $\hat{q}_{\delta}=\hat{q}-q_{0}$, where $q_{0}$ adds a constant offset. The physical charge operator $\hat{q}$ has the same commutation relation $\hat{\phi},\hat{q}]=i\hbar$ and so we can derive the same circuit dynamics (e.g. resonance frequency, nonlinearities, etc.) using the differential charge.
Three-wave mixing dynamics arise when charge on the capacitor is modulated at twice the resonator’s frequency. For example, consider a bias charge that includes both a DC component $q_{0}$ and an AC component $q_{\mathrm{drive}}(t)=q_{\mathrm{ac}}\cos(2\omega_{0}t+\theta)=q_{\mathrm{ac}}(e^{i(2\omega_{0}+\theta)}+e^{-i(2\omega_{0}+\theta)})/2$. The differential charge across the capacitor thus becomes $q_{\delta}\to q_{\delta}+q_{\mathrm{ac}}\cos(2\omega_{0}t+\theta)$, and the third-order term in Eq. 18 is expanded to $\frac{1}{3!}U_{c}^{(3)}(q_{0})(q_{\delta}+q_{\mathrm{drive}}(t))^{3}$.
Around the resonance frequency $\omega_{0}$ and within the rotating wave approximation (RWA), we are interested in the term $(q_{\delta}+q_{\mathrm{drive}}(t))^{3}\to 3q_{\delta}^{2}q_{\mathrm{drive}}(t)$ which will mix with the drive. We can re-write this term as a function of $a$ and $a^{\dagger}$, Eq. 19. In the RWA the terms $a^{2}$ and $a^{\dagger 2}$ remain because they rotate at $\pm 2\omega_{0}$, countering the drive, and the third order term thus becomes $U_{c}^{(3)}(q_{0})q_{\mathrm{zpf}}^{2}q_{\mathrm{ac}}(a^{2}e^{i\theta}+a^{\dagger^{2}}e^{-i\theta})/4$.
Under the RWA, the four-order (Kerr) term of the interaction Hamiltonian can be found by substituting $q_{\delta}$ into $\frac{1}{4!}U_{c}^{(4)}(q_{0})(a-a^{\dagger})^{4}\to\frac{1}{4!}U_{c}^{(4)}(q_{0})q_{\mathrm{zpf}}^{4}(6a^{\dagger 2}a^{2})$. In the quantum mechanical picture this expansion also renormalizes the resonance frequency to $\omega_{0}\to\omega_{0}+K_{\mathrm{eff}}$ due to the operator’s commutation relation; in our system this is a negligible effect because $K_{\mathrm{eff}}$ is small compared to the resonator’s linewidth. In the RWA, the Kerr term does not depend on the amplitude of a drive at $\pm 2\omega_{0}$. A drive near $\pm\omega_{0}$ would lead to four-wave mixing dynamics, but is not considered here.
In summary, we derive the effective Hamiltonian:
$H/\hbar=\omega_{0}a^{\dagger}a+\frac{\xi}{2}a^{\dagger 2}+\frac{\xi^{*}}{2}a^{2}+\frac{K_{\mathrm{eff}}}{2}a^{\dagger 2}a^{2}+\cdots.$ (20)
Here $\omega_{0}=\sqrt{U_{c}^{(2)}(q_{0})/L}=1/\sqrt{LC(v_{0})}$ is the resonator frequency, and $\xi$ and $K_{\mathrm{eff}}$ are the three-wave mixing (3WM) and effective Kerr terms, respectively:
$\xi(q_{0})=-\frac{U_{c}^{(3)}(q_{0})q_{\mathrm{ac}}q_{\mathrm{zpf}}^{2}}{2\hbar}e^{-i\theta},$ (21)
$K_{\mathrm{eff}}(q_{0})=\frac{U_{c}^{(4)}(q_{0})q_{\mathrm{zpf}}^{4}}{2\hbar}.$ (22)
Notice that $\omega_{0}$, $\xi$ and $K_{\mathrm{eff}}$ all depend on DC bias as given by derivatives of $U_{c}$ with respect to $q$, evaluated at $q_{0}$.
We can also express the three-wave mixing and effective Kerr terms as functions of the bias voltage $v_{0}$. This is a natural variable for use in experiments. To derive this dependence, we apply the product and quotient rules from calculus to the derivatives of $U_{c}(q)$, keeping in mind that $\partial U_{c}/\partial q=v(q)$, and $\partial^{2}U_{c}/\partial q^{2}=\partial v/\partial q=C^{-1}$. Doing so we get:
$\xi(v_{0})=\frac{C^{\prime}(v_{0})v_{\mathrm{ac}}v_{\mathrm{zpf}}^{2}}{2\hbar}e^{-i\theta},$ (23)
$K_{\mathrm{eff}}(v_{0})=\left(-C^{\prime\prime}(v_{0})+3\frac{C^{\prime}(v_{0})^{2}}{C(v_{0})}\right)\frac{v_{\mathrm{zpf}}^{4}}{2\hbar},$ (24)
where $C^{\prime}(v_{0})$ and $C^{\prime\prime}(v_{0})$ are the $1^{\mathrm{st}}$ and $2^{\mathrm{nd}}$ derivatives of capacitance with respect to voltage, respectively, evaluated at the bias voltage $v_{0}$. The zero-point voltage fluctuations are $v_{\mathrm{zpf}}=q_{\mathrm{zpf}}/C(v_{0})$ and, in Eq. 23, $v_{\mathrm{ac}}=q_{\mathrm{ac}}/C(v_{0})$ is the amplitude of the AC voltage drive across the capacitor at $2\omega_{0}$. We plot solutions to Eq. 23 and Eq. 24 in Fig. 2.
The magnitude of both of these nonlinearities depends on capacitor geometry, Fig. 5. In general, nonlinearities are greater for smaller mode volumes where the field is more concentrated. For example, a nanofabricated parallel plate capacitor with an order 100 nm separation made from STO or KTO is predicted to have 3WM terms of order MHz and an effective Kerr constant of order mHz; performing favorably as a three-wave mixing element in comparison to kinetic inductance parametric amplifiers (e.g. [21]). For larger, chip-scale dimensions (e.g. an order $100\,\mathrm{\mu m}$ gap), then the ratio of $\xi/K_{\mathrm{eff}}$ increases somewhat but the absolute value of both nonlinearities decreases. This is not ideal from the perspective of designing three-wave mixing elements, as both a large ratio of $\xi/K_{\mathrm{eff}}$ and a large absolute value of $\xi$ compared to other rates (e.g. loss) are desired so that strong parametric behavior can be achieved for small modulation amplitudes.
Figure 5: (a) The three-wave mixing (3WM) term $\xi/2\pi$ and (b) the effective Kerr term $K_{\mathrm{eff}}/2\pi$, both as functions of capacitor plate separation $d$, which is swept while keeping the ratio $A/d=(4\penalty 10000\ \,\mathrm{\mu m})^{2}/(200\penalty 10000\ \,\mathrm{nm})=8\times 10^{-5}\penalty 10000\ \,\mathrm{m}$ fixed. The 3WM term is reported at its maximum value, occuring at $v_{0}\neq 0$ and taking $v_{\mathrm{ac}}=1\,\mathrm{mV}$. The effective Kerr term is also reported at its maximum value, occuring $v_{0}=0$. Both nonlinearities decreases as the capacitor plates become further apart.
Appendix D Nanofabrication

### v2

@@ IV A Degenerate Parametric Amplifier
In Figs. 3d,e we plot $R(\omega)$ for both STO and KTO capacitors, choosing the dc bias voltage $v_{0}=v_{\mathrm{0,max}}$ to maximize $\xi$ for each design. For this operating point we simulate a center frequency and 3-dB bandwidth of $\omega_{0}/2\pi=2.072\,\mathrm{GHz}$ and $\kappa/2\pi=24\,\mathrm{MHz}$ for the STO design, respectively, and $\omega_{0}/2\pi=4.882\,\mathrm{GHz}$ and $\kappa/2\pi=49.5\,\mathrm{MHz}$ for the KTO design, respectively. The operating frequency can be tuned in situ by changing the dc bias. It can also be modified by designing different capacitor dimensions or a different inductance.
@@
The compression power of a degenerate parametric amplifier is a critical parameter that determines the maximum gain and operating range of the amplifier. It is set by a combination of effects including Kerr nonlinearity, higher-order terms in the circuit Hamiltonian, and pump depletion; as such it is complicated to precisely model. Inductive nonlinearities, e.g. kinetic inductance of the electrodes and the proposed stepped-impedance coupler, may also affect power compression. Though we present the existence of a Kerr-free operating point, we estimate from lowest-order theory, only, that a PANDA can have an effective Kerr term of order $K_{\mathrm{eff}}/2\pi\approx 0.1\,\mathrm{Hz}$, in comparison to a linewidth of $\kappa/2\pi\approx 20\,\mathrm{MHz}$ as shown in Fig. 3. It will therefore require $N\approx 10^{8}$ circulating photons to shift the resonator frequency by a linewidth. At a resonant frequency of 2 GHz, this corresponds to a circulating power of $P_{\mathrm{circ}}=N\hbar\omega_{0}\kappa\approx-64\,\mathrm{dBm}$, similar to state-of-the-art KIPAs (e.g. 1-dB compression powers of -50 dBm in Ref. [21], -65 dBm in Ref. [24]), and greater than JJ-based parametric amplifiers engineered for high-power compression (e.g. approximately -100 dBm in Ref. [6]). A three-wave-mixing element based on a nanoscale capacitor made from STO or KTO can therefore be expected to have power handling competitive with respect to superconducting alternatives, along with advantages from a compact form factor. To optimize power handling it is helpful to use thick electrodes to reduce kinetic inductance, and to use high-quality materials with large dielectric susceptibility and low loss.
V Conclusion
@@
We propose the development of a parametric mixing element based on the tunable permittivity of the quantum paraelectric perovskite materials STO and KTO. First, we derive an interaction Hamiltonian for a novel parametric device, realized by integrating this nonlinear element into a resonant circuit. We calculate voltage-tunable third- and fourth-order nonlinear terms that allow for three-wave- or four-wave-mixing dynamics, respectively, in analogy with superconducting parametric amplifiers based on the Josephson effect or kinetic inductance. We predict that the third-order (fourth-order) terms can have order MHz (sub-Hz) magnitude, a large ratio which suggests promise as a degenerate parametric amplifier (PANDA) with favorable dynamic range compared to superconducting alternatives. Due to the high permittivity of STO and KTO, a PANDA can also have a compact layout even at sub-GHz frequencies. We also note that with related materials or modifications from isotope exchange, strain, doping can lead to even stronger nonlinearities and lower loss near zero temperature than modeled here [47].
@@
While we focus specifically on a simple parametric amplifier circuit, we expect the large and tunable permittivity of STO and KTO to be widely useful in other rf and microwave frequency quantum circuits including filters, switches and modulators, nonreciprocal devices, and qubits. Unlike superconducting nonlinearities, STO and KTO are, in principle, robust to optical illumination and magnetic fields [55, 56]; PANDAs can thus be of use for applications such as microwave-to-optical transduction or axion dark-matter detection, where robust parametric mixing elements are desired. We predict application to next-generation axion dark-matter haloscopes will be especially fruitful; the combined need for tesla-scale magnetic fields, quantum-limited microwave measurement, and microwave squeezed light in modern experiments [17], remains a challenging operational limit for conventional superconducting parametric amplifiers. In summary, our proposal highlights the need for materials and nanofabrication development of STO and KTO microwave devices, and the unexplored application of these devices toward compact and scalable quantum technologies.
VI Acknowledgments
@@
The work in the Vučković group has been supported by the Department of Energy under the Q-NEXT program and by a Google AI Research grant. C.S.W. acknowledges the support of the Grainger Fellowship from the University of Chicago. G.S. acknowledges support from the Stanford Bloch Postdoctoral Fellowship. C.P.A is supported by AFOSR FA9550-25-1-0309. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. We thank Christian Haffner, Nathaniel Kinsey, Andrew Higginbotham, Osmond Wen and Nicholas Frattini for helpful discussions.
@@ Appendix A Dielectric Properties
The dielectric properties STO and KTO are well described by a modified Landau-Ginzburg-Devonshire (LGD) theory [41]. In this model, the permittivity is inversely proportional to the curvature of the Gibbs free energy as a function of atomic displacement [57, 58]. At room temperature, both STO and KTO are in a paraelectric phase, where the net atomic displacement is zero. Both materials approach a phase transition at cryogenic temperatures. However, this transition is suppressed by quantum fluctuations, resulting a quantum paraelectric state. In these systems, the free energy is described by two potential wells separated by a barrier, and which correspond to two directions of atomic displacement, Fig. 4a. At low temperatures, the energy barrier is weak enough compared to quantum fluctuations such that the atom occupies a superposition of both wells [37, 38].
Figure 4: At cryogenic temperature the perovskites STO and KTO have multiple low-energy states associated with positive or negative displacement of the central atom [40, 41]. This leads to a quantum paraelectric phase in which the quantum fluctuations (horizontal white arrows) are significant compared to the energy barrier between these two potential wells. Permittivity is proportional to the inverse curvature of the free energy compared to relative displacement. An electric field, $E$, tilts the potential, which constraints the magnitude of fluctuations and reduces permittivity. STO and KTO differ in the depth of the double-well potential and its susceptibility to an applied field [58]. (b) Relative permittivity $\varepsilon_{r}=\varepsilon/\varepsilon_{0}$ and (c) loss tangent $\mathrm{tan}(\delta)$, modeled vs $E$. Calculations use Eq. 5, which is based on the model and material parameters in Ref. [39].
Following work and notation by Vendik et al. [59], the dielectric constant of both STO and KTO are modeled using the modified LGD theory as
@@
describes the normalized biasing from the unpolarized state, where $E$ is the bias field, $E_{N}$ is a renormalizing field, and $\lambda_{S}$ is a function of defects/inhomogeneity. For a pure material with negligible defect density, then $\lambda_{S}\approx 0$ and $\lambda\to E/E_{N}$. Finally, the predicted frequency rolloff of the large dielectric response relates to the soft phonon frequency approxiately equal to THz, so at the relevant frequencies in the GHz range this assumption holds.
According to Ref. [39], the loss term $\Gamma(\lambda)=\Gamma_{1}(\lambda)+\Gamma_{2}(\lambda)+\Gamma_{3}$ in Eq. 5 is a combination of the following. (1) Loss related to multiphonon scattering, which scales with temperature such that $\Gamma_{1}(\lambda)=A_{1}(T/T_{c})^{2}G(\lambda)^{1/2}$, where $A_{1}$ is a constant. (2) Residual piezoelectricity, such that $\Gamma_{2}(\lambda)=A_{2}y(\lambda)$ where $A_{2}$ is a constant and
@@
is a residual ferroelectric displacement. Finally, (3) piezoelectricity due to charged defects such that $\Gamma_{3}=A_{3}n_{d}$, where $n_{d}$ is the density of charged defects and $A_{3}$ is another proportionality constant. The overall loss tangent is therefore modeled as
@@
These dielectric properties are relevant at rf and microwave frequency regimes and lower (e.g. characterized at 50 MHz in Ref. [42], 1 GHz in Ref. [45], 3 GHz in Ref. [60], and between 4 and 6 GHz in Ref. [48]). At frequencies of order THz and higher, the permittivity decreases due to oscillations associated with optical phonon modes in the crystal, and at optical frequencies STO has a refractive index of only 2.4 [47, 46]. To the author’s knowledge, the precise transition between these two limits has not yet been experimentally characterized but is predicted to be above 100’s of GHz [42, 43].
@@
Capacitance is proportional to permittivity, which in this case is voltage tunable.
@@ Appendix C Circuit Hamiltonian
where the capacitor’s energy is modified from the usual quadratic to account for a nonlinear capacitance (i.e. a nonlinear charge-voltage relation). This Hamiltonian is comprised of both a linear component $H_{\mathrm{harm}}$ and nonlinear interaction terms $H_{\mathrm{int}}$, such that $H=H_{\mathrm{harm}}+H_{\mathrm{int}}$.
@@
It is often helpful to express Eq. 18 in terms of the complex variables $a=\frac{1}{\sqrt{2\hbar}}\left(\frac{\phi}{\sqrt{z_{0}}}+i\sqrt{z_{0}}q_{\delta}\right)$ and $a^{\dagger}=\frac{1}{\sqrt{2\hbar}}\left(\frac{\phi}{\sqrt{z_{0}}}-i\sqrt{z_{0}}q_{\delta}\right)$ such that
@@
where $\phi_{\mathrm{zpf}}=\sqrt{\hbar z_{0}/2}$ and $q_{\mathrm{zpf}}=\sqrt{\hbar/2z_{0}}$ are the zero-point fluctuations of branch flux and charge, respectively [50]. Although for now we treat $\phi$ and $q_{\delta}$ as classical fields, in the quantum mechanical picture these become operators with the commutation relations $\hat{a},\hat{a}^{\dagger}]=1$ and $\hat{\phi},\hat{q}_{\delta}]=i\hbar$. In general these variables depend on bias charge since the resonator impedance is $z_{0}(v)=\sqrt{L/C(v)}$.
In order to analyze the underlying nonlinearity of our circuit, we first cast the total Hamiltonian in terms of the quadrature operators as follows:
$H/\hbar=\omega_{0}a^{\dagger}a+ig_{3}\big(a-a^{\dagger}\big)^{3}+g_{4}\big(a-a^{\dagger}\big)^{4}+$..., (21)
where $\omega_{0}(v)=1/\sqrt{LC(v)}$ is the resonance frequency, and $g_{3}$ and $g_{4}$ are the third- and fourth-order nonlinearities. It follows from Eq. 19 that
$g_{3}$ $=\frac{1}{3!}\frac{1}{\hbar}U^{(3)}_{c}q^{3}_{\mathrm{zpf}},$ (22)
$g_{4}$ $=\frac{1}{4!}\frac{1}{\hbar}U^{(4)}_{c}q^{4}_{\mathrm{zpf}}.$ (23)
The Taylor coefficients can be related to the external voltage bias $v_{0}$ keeping in mind that $\partial U_{c}/\partial q=v(q)$, and $\partial^{2}U_{c}/\partial q^{2}=\partial v/\partial q=C^{-1}$. One can show
$U^{(3)}_{c}(v_{0})$ $=-C^{\prime}(v_{0})/C^{3}(v_{0}),$ (24)
$U^{(4)}_{c}(v_{0})$ $=\bigg(-C^{\prime\prime}(v_{0})+3\frac{C^{\prime}(v_{0})^{2}}{C(v_{0})}\bigg)/C^{4}(v_{0}).$ (25)
Thus we can directly express our nonlinearities in terms of our capacitance:
$g_{3}(v_{0})$ $=\frac{1}{\hbar}\frac{v^{3}_{\mathrm{zpf}}C^{\prime}(v_{0})}{3!},$ (26)
$g_{4}(v_{0})$ $=\frac{1}{\hbar}\frac{v^{4}_{\mathrm{zpf}}}{4!}\left(-C^{\prime\prime}(v_{0})+3\frac{C^{\prime}(v_{0})^{2}}{C(v_{0})}\right),$ (27)
where $v_{\mathrm{zpf}}=q_{\mathrm{zpf}}/C(v_{0})$ are the resonator’s zero-point voltage fluctuations.
Three-wave mixing (also sometimes referred to as squeezing in the parametric amplifier literature) arises when charge on the capacitor is modulated at twice the resonator’s frequency. The typical approach to determine $\xi$ under the rotating-wave approximation (RWA) is to consider adding a modulation term to $q_{\delta}\to q_{\delta}+q_{\mathrm{ac}}\cos(2\omega_{0}t+\theta)$ and expanding Eq. 21 keeping resonant terms [51, 3]. This results in the Hamiltonian:
$H/\hbar=\omega_{0}a^{\dagger}a+\frac{\xi}{2}a^{\dagger 2}+\frac{\xi^{*}}{2}a^{2}+\frac{K_{\mathrm{eff}}}{2}a^{\dagger 2}a^{2}+\cdots,$ (28)
which appears in Eq. 3 of the main text. Here the three-wave-mixing term $\xi$ is related to Eq. 26 as
$\xi=3g_{3}\left(\frac{v_{\mathrm{ac}}}{v_{\mathrm{zpf}}}\right)e^{-i\theta},$ (29)
where $\theta$ is the phase of the pump tone and $v_{\mathrm{ac}}=q_{\mathrm{ac}}/C(v_{0})$ is the ac modulation voltage. To lowest order the effective Kerr term $K_{\mathrm{eff}}$ is related to Eq. 27 as
$K^{\mathrm{RWA}}_{\mathrm{eff}}=12g_{4}.$ (30)
One can also go further to second order in perturbation theory for Eq. 21 [52], which results in
$K^{\mathrm{2PT}}_{\mathrm{eff}}=12(g_{4}-5g^{2}_{3}/\omega_{0}).$ (31)
As shown in Fig. 3c of the main text, Eq. 31 therefore predicts an external bias where $K_{\mathrm{eff}}$ approaches zero.
Appendix D Design Considerations
In this section, we elucidate our circuit design considerations and attempt to separate out contributions from the PPC geometry and intrinsic material properties.
A critical design parameter is the STO or KTO film thickness $d$. In general, nonlinearities are greater for smaller mode volumes where the field is more concentrated. In Fig. 5, we analyze how the choice of $d$ impacts our circuit parameters while keeping the overall capacitance at zero external bias fixed (i.e. modifying $A$ and $d$ together such that $A/d$ remains constant). We observe that thinner capacitors result in higher nonlinearities, which translates to lower drive strengths required to obtain a target value of $\xi$, via Eq. 29. Understanding the trade-off between film thickness and internal loss will be key in optimizing performance.
Figure 5: Effect of varying the thin film thickness $d$ while keeping $A/d=(4\penalty 10000\ \,\mathrm{\mu m})^{2}/(200\penalty 10000\ \,\mathrm{nm})=8\times 10^{-5}\penalty 10000\ \,\mathrm{m}$ fixed for STO (top) and KTO (bottom). Thinner films have higher nonlinearities.
Operational bandwidth is another key design consideration. For a parametric amplifier based on a single resonant circuit this is set by the resonator’s linewidth, $\kappa=\kappa_{\mathrm{ext}}+\kappa_{\mathrm{int}}$. Linewidth is a combination of external coupling rate $\kappa_{\mathrm{ext}}$ that may be near arbitrarily determined by choice of coupler, and, internal loss, which here we model to be dominated by the dielectric loss tangent $\kappa_{\mathrm{int}}=\omega_{0}\tan(\delta)$. In Fig. 6 we plot $\kappa_{\mathrm{int}}$ versus resonator frequency, for the dielectric model based on LGD theory as described in Appendix A. In general, it is best to design a single-port parametric amplifier to be overcoupled, i.e. $\kappa_{\mathrm{ext}}\gg\kappa_{\mathrm{int}}$, because internal loss scales added noise by a factor of $(\kappa_{\mathrm{int}}+\kappa_{\mathrm{ext}})/\kappa_{\mathrm{ext}}$ [51]. For the parameters described here, chosen around an operation frequency at approximately $\mathrm{GHz}$ frequencies, this suggests that an external coupling rate of 10’s of MHz or greater will yield the best noise performance for both PANDA designs.
The resonator’s internal and external loss are related to its gain-bandwidth by solutions to Eq. 4. To maximize linear gain it is useful to pump the circuit close to $2\xi/\kappa=1$, the limit of parametric instability, as plotted in Fig. 3. The aforementioned constraints on $\xi$ thus determine that $\kappa_{\mathrm{ext}}$ cannot be arbitrarily increased without compromise to gain. As such, we predict in Fig. 3 that a gain of greater than 20 dB over order MHz of bandwidth is reasonable given design constraints. More bandwidth may require a more complicated circuit, paralleling work on broadband superconducting parametric amplifiers [5, 20, 22].
Figure 6: Internal loss rate $\kappa_{\mathrm{int}}=\omega_{0}\times\mathrm{tan}(\delta)$ vs. resonator frequency, for both devices, where the loss tangent is determined from Landau-Ginzburg-Devonshire theory (see Fig. 4c).
Appendix E Nanofabrication